OpenStudy (anonymous):

Can someone PLEASE tell me if I'm right? A computer company uses the function f(h) = 50h + 75 to calculate the total repair bill, f(h), for h number of hours worked. Find h if f(h) = $275. Is the answer 4 hours?

Vocaloid (vocaloid):

yes, 4 is the correct answer 275 = 50(h) + 75 200 = 50(h) h=4

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (anonymous):

I also need another question checked. The function d(x) = 3x + 2 determines how many donuts need to be purchased for a meeting, where x is the number of people at the meeting. The office manager uses h(d(x)) to find the amount of money to bring for the donut purchase. The function h(x) = 2x + 7. Solve for how much money to bring when there are 8 people in the meeting. Is the answer 59?

Vocaloid (vocaloid):

yes, 59 is correct d(8) = 3(8) + 2 = 26 h(26) = 2(26) + 7 = 59

OpenStudy (anonymous):

thank you :)